two spherical objects have masses of 7.5x10^3 kg and 9.2x19^7 kg. Their centers are separated by a distance of 2.5 x 10^3m. Find the gravitational attraction between them


1
Expert's answer
2021-06-03T18:25:45-0400

The gravitational attraction can be found by the law of universal gravitation:


"F=G\frac{m_1m_2}{R^2},\\\space\\
(6.673·10^{-11})\frac{(7.5·10^3) (9.2·19^7)}{(2.5·10^3)^2}=7.37·10^{-6}\text{ N}."